OpenStudy (anonymous):

3^x = 3^kx find k hence find int 3^x dx

OpenStudy (anonymous):

what are we supposed to do? find k? then k*ln3?

OpenStudy (anonymous):

after find k and find int 3^x dx

OpenStudy (anonymous):

to me, k =1, since we have \(3^x =3^{kx}\rightarrow ~~x =kx \) iff k =1

OpenStudy (anonymous):

sorry k is a constant

OpenStudy (anonymous):

no difference. k =1 and 1 is a constant, too. for the second part \(\int 3^x dx =3^x +C\)

OpenStudy (anonymous):

thanks